solve each of the following sysytems by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
2x + 3y =1 and 5x + 3y =16

Answer by ankor@dixie-net.com(22740)

2x + 3y = 1
5x + 3y = 16
:
Multiply the 1st equation by -1 and add:
-2x - 3y = -1
+5x + 3y = 16
-------------- adding eliminates y, solve for x
3x = 15
x = 15/3
x = 5
:
Find y using 2x + 3y = 1, substitute 5 for x:
2(5) + 3y = 1
10 + 3y = 1
3y = 1 - 10
3y = -9
y = -9/3
y = -3
:
:
Check solutions in 5x + 3y = 16, substitute for x and y:
5(5) + 3(-3) =
25 - 9 = 16